OpenStudy (anonymous):

Solve for x and y. 2x + y=1 6x − 3y=5

OpenStudy (rulnick):

easiest: triple the first equation, then add the equations

OpenStudy (rulnick):

... after tripling, first eqn becomes 6x+3y=3

OpenStudy (anonymous):

x=2/3

OpenStudy (anonymous):

y= -1/3

OpenStudy (texaschic101):

2x + y = 1 y = -2x + 1 now sub -2x + 1 in for y in the other equation 6x - 3y = 5 6x - 3(-2x + 1) = 5 6x + 6x - 3 = 5 6x + 6x = 5 + 3 12x = 8 x = 8/12 x = 2/3 now sub 2/3 in for x in either equation to find y 2x + y = 1 2(2/3) + y = 1 4/3 + y = 1 y = 1 - 4/3 y = 3/3 - 4/3 y = -1/3 check.. 6x - 3y = 5 2x + y = 1 6(2/3) - 3(-1/3) = 5 2(2/3) + (-1/3) = 1 4 + 1 = 5 4/3 - 1/3 = 1 5 = 5 (correct) 3/3 = 1 1 = 1 (correct) x = 2/3 and y = -1/3 any questions ?

OpenStudy (anonymous):

@texaschic101 A well deserved medal

OpenStudy (texaschic101):

thank you :)

OpenStudy (anonymous):

Thx for the help